# Online Decision-Focused Learning

### v2

@@ Abstract
Decision-focused learning (DFL) is an increasingly popular paradigm for training predictive models whose outputs are used in decision-making tasks. Instead of merely optimizing for predictive accuracy, DFL trains models to directly minimize the loss associated with downstream decisions. However, existing studies focus solely on scenarios where a fixed batch of data is available and the objective function does not change over time. We instead investigate DFL in dynamic environments where the objective function and data distribution evolve over time. This setting is challenging for online learning because the objective function has zero or undefined gradients—which prevents the use of standard first-order optimization methods—and is generally non-convex. To address these difficulties, we (i) regularize the objective to make it differentiable and (ii) use perturbation techniques along with a near-optimal oracle to overcome non-convexity. Combining those techniques yields two original online algorithms tailored for DFL, for which we establish respectively static and dynamic regret bounds. These are the first provable guarantees for the online decision-focused problem. Finally, we showcase the effectiveness of our algorithms on a knapsack experiment, where they outperform two standard benchmarks.
1 Introduction.
@@
Many real-world decision problems involve uncertainty, and a common approach to handling it is through the predict-then-optimize framework [10]. First, a prediction model is trained on historical data; then, its output is fed into an optimization problem to guide decision-making. This natural strategy has been successfully applied in many operation research (OR) problems, ranging from supply chain management [2, 28, 7, 10] and revenue management [26, 27, 18, 16] to healthcare operation [11, 5, 34, 61], see Mivsić and Perakis [56] for an extensive review. It is clear that this approach would yield optimal decisions if the predictions were perfectly accurate. However, in practice, prediction errors are inevitable—and even small inaccuracies can propagate through the optimization process, potentially leading to poor decisions.
@@
To address this limitation, an approach known as decision-focused learning [54]—also referred to as smart predict-then-optimize [25] or integrated learning-optimization [62]—has emerged. Instead of optimizing for prediction accuracy alone, this method trains the predictive model to directly minimize the downstream decision loss. By aligning the learning objective with the decision-making goal, it produces models that are more robust to prediction errors in practical applications. While decision-focused learning yields strong empirical performance [23, 71, 72, 73], theoretical development has so far been limited to the batch setting, where models are trained on pre-collected, independently and identically distributed (i.i.d.) data [74, 53, 64, 63]. This assumption breaks down in many real-world scenarios involving dynamic environments [17, 57] and shifting data distributions [50, 60].
@@
Online learning [15, 37] provides a general way to cope with such non-stationarity of data-generating processes. This framework considers a learner who makes decisions sequentially, at each round leveraging data collected from previous rounds to inform its next decision. Crucially, the objective function is allowed to vary over time, either in a stochastic or adversarial way. This provides a natural framework for the work that we present here, which extends decision-focused learning beyond the i.i.d., batch setting.
@@
We develop a theoretical foundation for online decision-focused learning, enabling its application in non-stationary settings. This presents a significant technical challenge, as the inherent difficulties of decision-focused learning—such as the non-differentiability of the objective function or the lack of convexity of the losses due to the bi-level nature of the problem—compound those already present in online learning. Our contributions are as follows:
@@
    We present two algorithms to tackle this problem, Decision-Focused Follow-the-Perturbed-Leader (DF-FTPL) and Decision-Focused Online Gradient Descent (DF-OGD). While both rely on regularizing the inner problem to make the resulting decision differentiable, they differ on the way they update the parameters of the prediction model. DF-FTPL uses the FTPL approach [43], and DF-OGD leverages a variant of Online Gradient Descent [81]. We establish sublinear convergence guarantees for both procedures, in the form of a static regret bound for the former and a dynamic regret bound for the latter. To our knowledge, these are the first provable guarantees for the online decision-focused learning problem.
@@
    Finally, we assess the performance of our algorithms on a knapsack experiment inspired by [54]. Our simulations demonstrate that our approach outperforms the online version of two popular baselines, namely prediction-focused learning and Smart-Predict-then-Optimize, in both static and dynamic environments.
@@
Our work addresses the decision-focused learning problem, pioneered by [25]. Several approaches have been proposed in the batch setting, among which differentiating the associated KKT conditions [33, 4, 23, 74, 52], smoothing via random perturbation [9], building surrogate losses via duality [25] or directional gradients [42] and relying on pairwise ranking techniques [53], see [54, 62] for additional references. However, the extension of decision-focused learning to the online setting remains unexplored. While several recent studies address online bi-level optimization [65, 69, 49], none of them are applicable to decision-focused learning as they rely on restrictive smoothness assumptions, that are incompatible with the structure of decision-focused problems. In particular, the decision-focused objective is typically non-convex and features gradients that are either zero or undefined, owing to its underlying linear structure.
This motivates the use of different online methods, specifically tailored to address these problems. On the one hand, lack of differentiability is usually tackled through zero-th order methods [39, 29], sub-gradient [24], proximal [22] or smoothing approaches [1]. On the other hand, methods to address non-convexity in online learning often rely on the existence of a near optimal oracle [45, 3, 67, 75] or additional smoothness conditions [46, 32]. All those results are derived for static regret [81] which compares the learned predictors to the best static strategy and is defined in Section˜2. A more challenging criterion, also described in Section˜2 is dynamic regret, introduced in [81] and later developed in [35, 12, 79, 78] among others, which takes into account the evolution of the environment. An alternative approach to obtain guarantees in the online non-convex setting is to weaken static regret to local regret [38, 6, 80, 36], which is the sum of the objective gradient norms evaluated in the iterates over time. Minimizing local regret corresponds to encouraging convergence toward stationary points. However as we shall see, this approach is not sensible in our framework, as objective gradients are either zero or undefined.
@@
For a differentiable map $\varphi:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that $\mathcal{Y}\subseteq\mathbb{R}$, $\nabla\varphi(x)$ denotes the gradient of $g$ at $x\in\mathbb{R}^{m}$ and $\nabla^{2}\varphi(x)$ denotes the Hessian of $g$. In the case where $\mathcal{Y}\subseteq\mathbb{R}^{d}$, $\nabla\varphi(x)$ denotes the Jacobian of $\varphi$. For two vectors $(v,w)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $v\succcurlyeq w$ means that $v_{i}\geqslant w_{i}$ for any $i\in[d]$, and $\mathopen{}\mathclose{{\left\langle v,w}}\right\rangle=v^{\mathsf{\scriptscriptstyle T}}w$ refers to the standard Euclidian inner product. Also, $\mathopen{}\mathclose{{\left\|v}}\right\|=\sqrt{\mathopen{}\mathclose{{\left\langle v,v}}\right\rangle}$ is the standard euclidian norm. Given a compact convex set $\Theta\subseteq\mathbb{R}^{d}$, $\Pi_{\Theta}$ denotes the orthogonal projection onto $\Theta$. For a matrix $M\in\mathbb{R}^{m\times d}$, $\mathopen{}\mathclose{{\left\|M}}\right\|_{\mathrm{op}}$ refers to its $\mathrm{L}^{2}$-operator norm. In the case $d=m$, $\lambda_{\min}(M)$ refers to its lowest real eigenvalue and $\lambda_{\max}(M)$ it largest real eigenvalue. For $(x,y)\in\mathbb{R}^{2}$, $x\propto y$ means that there exists $\lambda\in\mathbb{R}^{\star}$ such that $x=\lambda y$.
@@ 2 Framework
where $X_{t}\in\mathcal{X}$ are covariates (possibly random) and $\mathcal{W}=\mathrm{Conv}(v_{1},\ldots,v_{K})$ is a bounded convex polytope of $\mathbb{R}^{d}$ with non-empty interior and vertices $\{v_{i}\}_{i=1}^{K}$. This feasible set appears naturally in many problems such as shortest-path [30], portfolio selection [47] or mixed strategy design in games [68]. In (1), $\bar{g}_{t}:\mathcal{X}\to\mathbb{R}^{d}$, is a cost function, which satisfies for instance $\bar{g}_{t}(X_{t})=\mathbb{E}[Z_{t}\mid X_{t}]$ for some hidden state $Z_{t}\in\mathbb{R}^{d}$. Importantly, $X_{t}$ is revealed at the beginning of the round, but not $\bar{g}_{t}(X_{t})$ which is only available at the end of the round. While $\bar{g}_{t}(X_{t})$ is unknown to the decision-maker at the decision time, they have access to a family of models $g:\Theta\times\mathcal{X}\rightarrow\mathbb{R}^{d}$, parameterized by $\Theta\subset\mathbb{R}^{m}$, to predict it. Then, the general form of the decision-making dynamics we consider can be described as follows: at each round $t\in[T]$, for a horizon $T\in\mathbb{N}$, given the current parameter $\theta_{t}$,
@@
  3. 3.
    Finally, they observe $\bar{g}_{t}(X_{t})$ and update $\theta_{t+1}$ for the next round.
Formally, the decision-maker considers a joint process $(\theta_{t},w_{t})_{t\in[T]}$, starting from $\theta_{1}\in\Theta$ and a history $\mathcal{H}_{0}=\emptyset$, and defined by the following recursion. At round $t\in[T]$, the decision-maker takes the best action given the collected history $\mathcal{H}_{t-1}$, the current estimate $\theta_{t}$ and a new feature $X_{t}$ as in (2). Then, they observe $\bar{g}_{t}(X_{t})$ and update their history $\mathcal{H}_{t}=\mathcal{H}_{t-1}\cup\{(X_{t},\bar{g}_{t}(X_{t}),\theta_{t},w^{\star}_{t}(\theta_{t}))\}$. Finally, they update their prediction parameter $\theta_{t+1}$ based on $\mathcal{H}_{t}$ through an algorithm $\mathsf{Alg}_{t}$. A central question for the decision-maker is then the choice of algorithms $\{\mathsf{Alg}_{t}\}_{t\in[T]}$ to fit their regression model, that is how to pick $\theta_{t}\in\Theta$ for each $t\in[T]$.
@@
From an online perspective, prediction-focused learning consists in selecting at each $t\in[T]$ an algorithm $\mathsf{Alg}_{t}$ to estimate $\operatorname*{\arg\!\min}_{\theta\in\Theta}\mathrm{R}_{t}(\theta)$, where $\mathrm{R}_{t}$ is a statistical risk based on the historical observations $\mathcal{H}_{t-1}$, typically chosen as the empirical risk $\mathrm{R}_{t}(\theta)=\sum_{i=1}^{t}\ell(\bar{g}_{i}(X_{i}),g(\theta,X_{i}))$, for some loss function $\ell$ (e.g., cross-entropy or squared error).
Alternatively, we consider in this paper, the decision-focused learning approach, which selects parameters $\theta\in\Theta$ by directly minimizing the downstream objective instead of a risk function. Specifically, the decision-maker chooses $\mathsf{Alg}_{t}$ to solve $\operatorname*{\arg\!\min}_{\theta\in\Theta}\langle\bar{g}_{t}(X_{t}),w_{t}^{\star}(\theta)\rangle$ where $w_{t}^{\star}$ is defined in (2). From the previous formulation, we see that the decision-focused learning formulation corresponds to a bilevel optimization problem [19, 66, 44].
We emphasize that no stationary assumptions are made on the process $\{X_{t},\bar{g}_{t}(X_{t})\}_{t\in[T]}$, allowing an adversarial selection of the features $X_{t}$ and cost functions $\bar{g}_{t}$. To fully accommodate this flexibility, we assess the optimality of $\{\theta_{t}\}_{t\in[T]}$ with two notions of regret. Let
@@
denotes the loss incurred when taking an action based on the prediction parameter $\theta_{t}\in\Theta$. Inspired from [81], we consider the notions of static and dynamic regret to measure the effectiveness of a learning strategy $\mathopen{}\mathclose{{\left\{\theta_{t}}}\right\}_{t\in[T]}$, which are respectively:
@@
where $\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}}\right]$. Note that our static regret compare to the best fixed strategy, with respect to the averaged losses. This notion makes sense here as we aim to control those regrets in expectation over the randomness of the process. In our dynamic regret, we consider the expectation of the loss given past information to prevent the dynamic comparator from overfitting the current realization of $X_{t}$.
@@
(i) $\Theta\subset\mathbb{R}^{m}$ is a compact and convex set with diameter $D_{\Theta}<\infty$. (ii) For any $\theta\in\Theta$ and $t\in[T]$, $\nabla_{\theta}g(\theta,X_{t})\in\mathbb{R}^{m\times d}$ is continuously differentiable and $\mathopen{}\mathclose{{\left\|\nabla_{\theta}g(\theta,X_{t})}}\right\|_{\mathrm{op}}\leqslant G<\infty$ almost-surely. (iii) For any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|\leqslant D_{\mathcal{Z}}<\infty$ almost-surely.
It is common in online learning to assume boundedness of the parameter space, model gradient and prediction space [14, 13]. Note that in the well-specified setting, i.e., $\bar{g}_{t}=g(\bar{\theta}_{t},\cdot)$ for some $\bar{\theta}_{t}$, these two former conditions automatically imply the latter. We emphasize however that we do not assume this realizability condition, in contrast to most of the literature on decision-focused learning [8].
Without further assumptions on the sequence of costs $\bar{g}_{t}$ and models $g$, the problem can still be made arbitrarily hard. Therefore, we make an assumption coming from the classification [51, 70] and bandit literature [76, 59] about the margins of the cost function. Recall that we denote by $\{v_{i}\}_{i=1}^{K}$ the vertices of $\mathcal{W}$. We define for any $x\in\mathcal{X}$, $u_{i}(\theta,x)=\mathopen{}\mathclose{{\left\langle g(\theta,x),v_{i}}}\right\rangle$ for $i\in[d]$.
@@
There exist $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{\left[0,1}}\right]$, such that almost surely for any $t\in[T]$, $\theta\in\Theta$ and $\varepsilon\in\mathopen{}\mathclose{{\left[0,1}}\right]$,
@@
3 Online algorithms for decision-focused learning .

### v3

@@ Abstract
Decision-focused learning (DFL) is an increasingly popular paradigm for training predictive models whose outputs are used in decision-making tasks. Instead of merely optimizing for predictive accuracy, DFL trains models to directly minimize the loss associated with downstream decisions. However, existing studies focus solely on scenarios where a fixed batch of data is available and the objective function does not change over time. We instead investigate DFL in dynamic environments where the objective function and data distribution evolve over time. This setting is challenging for online learning because the objective function has zero or undefined gradients, which prevents the use of standard first-order optimization methods, and is generally non-convex. To address these difficulties, we (i) regularize the objective to make it differentiable and (ii) use perturbation techniques along with a near-optimal oracle to overcome non-convexity. Combining those techniques yields two original online algorithms tailored for DFL, for which we establish respectively static and dynamic regret bounds. These are the first provable guarantees for the online decision-focused problem. Finally, we showcase the effectiveness of our algorithms on a knapsack experiment, where they outperform two standard benchmarks.
1 Introduction
@@
Many real-world decision problems involve uncertainty, which is commonly handled through the predict-then-optimize framework (Bertsimas and Kallus, 2020). First, a prediction model is trained on historical data; then, its output is fed into an optimization problem to guide decision-making. This natural strategy has been successfully applied in many operation research (OR) problems, ranging from supply chain management (Acimovic and Graves, 2015; Fisher et al., 2016; Ban and Rudin, 2019; Bertsimas and Kallus, 2020), revenue management (Farias et al., 2013; Ferreira et al., 2016; Cohen et al., 2017; Chen et al., 2022) to healthcare operation (Bertsimas et al., 2013; Aswani et al., 2019; Gupta et al., 2020; Rath et al., 2017), see Mivsić and Perakis (2020) for an extensive review. It is clear that this approach would yield optimal decisions if the predictions were perfectly accurate. However, in practice, prediction errors are inevitable, and even small inaccuracies can propagate through the optimization process, potentially leading to poor decisions.
@@
To address this limitation, an approach known as decision-focused learning (Mandi et al., 2024), also referred to as smart predict-then-optimize (Elmachtoub and Grigas, 2022) or integrated learning-optimization (Sadana et al., 2025), has emerged. Instead of optimizing for prediction accuracy alone, this method trains the predictive model to directly minimize the downstream decision loss. By aligning the learning objective with the decision-making goal, it produces models that are more robust to prediction errors in practical applications. While decision-focused learning yields strong empirical performance (Donti et al., 2017; Verma et al., 2022; 2023; Wang et al., 2023), theoretical development has so far been limited to the batch setting, where models are trained on pre-collected, independently and identically distributed (i.i.d.) data (Wilder et al., 2019; Mandi et al., 2022; Shah et al., 2022; Schutte et al., 2024). This assumption breaks down in many real-world scenarios involving dynamic environments (Cheung et al., 2019; Padakandla et al., 2020) and shifting data distributions (Lu et al., 2018; Quiñonero-Candela et al., 2022).
@@
Online learning (Cesa-Bianchi and Lugosi, 2006; Hazan, 2023) provides a general way to cope with such non-stationarity of data-generating processes. This framework considers a learner who makes decisions sequentially, at each round leveraging data collected from previous rounds to inform its next decision. Crucially, the objective function is allowed to vary over time, either in a stochastic or adversarial way. This provides a natural framework for the work that we present here, which extends decision-focused learning beyond the i.i.d., batch setting.
@@
We develop a theoretical foundation for online decision-focused learning, enabling its application in non-stationary settings. This presents a significant technical challenge, as the inherent difficulties of decision-focused learning, such as the non-differentiability of the objective function or the lack of convexity of the losses due to the bi-level nature of the problem, compound those already present in online learning. Our contributions are as follows:
@@
    We present two algorithms to tackle this problem, Decision-Focused Follow-the-Perturbed-Leader (DF-FTPL) and Decision-Focused Online Gradient Descent (DF-OGD). While both rely on regularizing the inner problem to make the resulting decision differentiable, they differ on the way they update the parameters of the prediction model. DF-FTPL uses the FTPL approach (Hutter and Poland, 2005), and DF-OGD leverages a variant of Online Gradient Descent (Zinkevich, 2003). We establish sublinear convergence guarantees for both procedures, in the form of a static regret bound for the former and a dynamic regret bound for the latter. To our knowledge, these are the first provable guarantees for the online decision-focused learning problem.
@@
    Finally, we assess the performance of our algorithms on a knapsack experiment inspired by Mandi et al. (2024). Our simulations demonstrate that our approach outperforms the online version of two popular baselines, namely prediction-focused learning and Smart-Predict-then-Optimize, in both static and dynamic environments.
@@
Several decision-focused approaches have been proposed in the batch setting, among which differentiating the associated KKT conditions Gould et al. (2016); Amos and Kolter (2017); Donti et al. (2017); Wilder et al. (2019); Mandi and Guns (2020), smoothing via random perturbation Berthet et al. (2020), building surrogate losses via duality Elmachtoub and Grigas (2022) or directional gradients Huang and Gupta (2024) and relying on pairwise ranking techniques Mandi et al. (2022), see Mandi et al. (2024); Sadana et al. (2025) for additional references. However, the extension of decision-focused learning to the online setting remains unexplored. While several recent studies address online bi-level optimization Shen et al. (2023); Tarzanagh et al. (2024); Lin et al. (2023), none of them are applicable to decision-focused learning as they rely on restrictive smoothness assumptions, that are incompatible with the structure of decision-focused problems. In particular, the decision-focused objective is typically non-convex and features gradients that are either zero or undefined, owing to its underlying linear structure.
This motivates the use of different online methods, specifically tailored to address these problems. On the one hand, lack of differentiability is usually tackled through zero-th order methods Héliou et al. (2020); Frezat et al. (2023), sub-gradient Duchi et al. (2011), proximal Dixit et al. (2019) or smoothing approaches Abernethy et al. (2014). On the other hand, methods to address non-convexity in online learning often rely on the existence of a near optimal oracle Kalai and Vempala (2005); Agarwal et al. (2019); Suggala and Netrapalli (2020); Xu and Zhang (2024) or additional smoothness conditions Lesage-Landry et al. (2020); Ghai et al. (2022). All those results are derived for static regret (Zinkevich, 2003) which compares the learned predictors to the best static strategy and is defined in Section˜2. A more challenging criterion, also described in Section˜2 is dynamic regret, introduced in Zinkevich (2003) and later developed in Hall and Willett (2013); Besbes et al. (2015); Zhao et al. ; Zhao and Zhang (2021) among others, which takes into account the evolution of the environment. An alternative approach to obtain guarantees in the online non-convex setting is to weaken static regret to local regret Hazan et al. (2017); Aydore et al. (2019); Zhuang et al. (2020); Hallak et al. (2021), which is the sum of the objective gradient norms evaluated in the iterates over time. Minimizing local regret corresponds to encouraging convergence toward stationary points. However as we shall see, this approach is not sensible in our framework, as objective gradients are either zero or undefined.
@@
For a differentiable map $\varphi:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that $\mathcal{Y}\subseteq\mathbb{R}$, $\nabla\varphi(x)$ denotes the gradient of $\varphi$ at $x\in\mathbb{R}^{m}$ and $\nabla^{2}\varphi(x)$ denotes the Hessian of $\varphi$. In the case where $\mathcal{Y}\subseteq\mathbb{R}^{d}$, $\nabla\varphi(x)$ denotes the Jacobian of $\varphi$. For two vectors $(v,w)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $v\succcurlyeq w$ means that $v_{i}\geqslant w_{i}$ for any $i\in[d]$, and $\mathopen{}\mathclose{{\left\langle v,w}}\right\rangle=v^{\mathsf{\scriptscriptstyle T}}w$ refers to the standard Euclidian inner product. Also, $\mathopen{}\mathclose{{\left\|v}}\right\|=\sqrt{\mathopen{}\mathclose{{\left\langle v,v}}\right\rangle}$ is the standard euclidian norm. Given a compact convex set $\Theta\subseteq\mathbb{R}^{d}$, $\Pi_{\Theta}$ denotes the orthogonal projection onto $\Theta$. For a matrix $M\in\mathbb{R}^{m\times d}$, $\mathopen{}\mathclose{{\left\|M}}\right\|_{\mathrm{op}}$ refers to its $\mathrm{L}^{2}$-operator norm. In the case $d=m$, $\lambda_{\min}(M)$ refers to its lowest real eigenvalue and $\lambda_{\max}(M)$ it largest real eigenvalue. For $(x,y)\in\mathbb{R}^{2}$, $x\propto y$ means that there exists $\lambda\in\mathbb{R}^{\star}$ such that $x=\lambda y$.
@@ 2 Framework
where $\mathcal{W}=\mathrm{Conv}(v_{1},\ldots,v_{K})$ is a bounded convex polytope of $\mathbb{R}^{d}$ with non-empty interior and vertices $\{v_{i}\}_{i=1}^{K}$. This feasible set appears naturally in many problems such as shortest-path (Gallo and Pallottino, 1988), portfolio selection (Li and Hoi, 2014) or mixed strategy design in games (Syrgkanis et al., 2015). In (1), $X_{t}$ are random covariates and $\bar{g}_{t}:\mathcal{X}\to\mathbb{R}^{d}$, is a deterministic cost function, which satisfies for instance $\bar{g}_{t}(X_{t})=\mathbb{E}[Z_{t}\mid X_{t}]$ for some hidden state $Z_{t}\in\mathbb{R}^{d}$. At each period $t$, nature picks both a distribution for $X_{t}$ and a cost function $g_{t}$. This situation corresponds to the stochastic adversary setting (Rakhlin et al., 2011). Importantly, $X_{t}$ is revealed at the beginning of the round, but not $\bar{g}_{t}(X_{t})$ which is only available at the end of the round. While $\bar{g}_{t}(X_{t})$ is unknown to the decision-maker at the decision time, they have access to a family of models $g:\Theta\times\mathcal{X}\rightarrow\mathbb{R}^{d}$, parameterized by $\Theta\subset\mathbb{R}^{m}$, to predict it. The general form of the decision-making dynamics we consider can be described as follows: at each round $t\in[T]$, for a horizon $T\in\mathbb{N}$, given the current parameter $\theta_{t}$,
@@
    Nature picks a distribution for $X_{t}\in\mathcal{X}$ and a cost function $\bar{g}_{t}:\mathcal{X}\rightarrow\mathcal{Z}$.
  2. 2.
@@
  4. 4.
    Finally, the decision-maker observes $\bar{g}_{t}(X_{t})$ and update $\theta_{t+1}$ for the next round.
Formally, the decision-maker considers a joint process $(\theta_{t},w_{t})_{t\in[T]}$, starting from $\theta_{1}\in\Theta$ and a history $\mathcal{H}_{0}=\emptyset$, and defined by the following recursion. At round $t\in[T]$, the decision-maker takes the best action given the current estimate $\theta_{t}$ and a new feature $X_{t}$ as in (2). Then, they observe $\bar{g}_{t}(X_{t})$ and update their history $\mathcal{H}_{t}=\mathcal{H}_{t-1}\cup\{(X_{t},\bar{g}_{t}(X_{t}),\theta_{t},w^{\star}_{t}(\theta_{t}))\}$. Finally, they update their prediction parameter $\theta_{t+1}$ based on $\mathcal{H}_{t}$ through an algorithm $\mathsf{Alg}_{t}$. A central question for the decision-maker is then the choice of algorithms $\{\mathsf{Alg}_{t}\}_{t\in[T]}$ to fit their regression model, that is how to pick $\theta_{t}\in\Theta$ for each $t\in[T]$.
@@
From an online perspective, prediction-focused learning consists in selecting at each $t\in[T]$ an algorithm $\mathsf{Alg}_{t}$ to estimate $\operatorname*{\arg\!\min}_{\theta\in\Theta}\mathrm{R}_{t}(\theta)$, where $\mathrm{R}_{t}$ is a statistical risk based on the historical observations $\mathcal{H}_{t}$, typically chosen as the empirical risk $\mathrm{R}_{t}(\theta)=\sum_{i=1}^{t}\ell(\bar{g}_{i}(X_{i}),g(\theta,X_{i}))$, for some loss function $\ell$ (e.g., cross-entropy or squared error).
Alternatively, we consider in this paper, the decision-focused learning approach, which selects parameters $\theta\in\Theta$ by directly minimizing the downstream objective instead of a risk function. Specifically, the decision-maker chooses $\mathsf{Alg}_{t}$ to solve $\operatorname*{\arg\!\min}_{\theta\in\Theta}\langle\bar{g}_{t}(X_{t}),w_{t}^{\star}(\theta)\rangle$ where $w_{t}^{\star}$ is defined in (2). From the previous formulation, we see that the decision-focused learning formulation corresponds to a bilevel optimization problem (Colson et al., 2007; Sinha et al., 2017; Ji et al., 2021).
We emphasize that no stationary assumptions are made on the process $\{X_{t},\bar{g}_{t}(X_{t})\}_{t\in[T]}$. In other words, the adversary may select any distribution for $X_{t}$ and any function $\bar{g}_{t}$ (as long as H˜2 below is satisfied). To fully accommodate this flexibility, we assess the optimality of $\{\theta_{t}\}_{t\in[T]}$ with two notions of regret. Let
@@
denotes the loss incurred when taking an action based on the prediction parameter $\theta_{t}\in\Theta$. Following Zinkevich (2003), we consider the notions of static and dynamic regret to measure the effectiveness of a learning strategy $\mathopen{}\mathclose{{\left\{\theta_{t}}}\right\}_{t\in[T]}$, which are respectively:
@@
where $\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}}\right]$. In our dynamic regret, the sequence of actions is compared against a sequence of oracles, each minimizing the instantaneous loss. Without taking the conditional expectation over $\mathcal{H}_{t-1}$, each comparator could overfit to the specific realization of $X_{t}$ resulting in an unrealistically strong benchmark. By considering the conditional expectation of the loss, we effectively regularize the dynamic comparators, making them meaningful competitors. Note that our static regret compare to the best fixed strategy, with respect to the averaged losses. This notion makes sense here as we aim to control those regrets in expectation over the randomness of the process.
@@
(i) $\Theta\subset\mathbb{R}^{m}$ is a compact and convex set with diameter $D_{\Theta}<\infty$. (ii) for any $t\in[T]$, $\theta\mapsto g(\theta,X_{t})\in\mathbb{R}^{m\times d}$ is continuously differentiable almost surely and $\mathopen{}\mathclose{{\left\|\nabla_{\theta}g(\theta,X_{t})}}\right\|_{\mathrm{op}}\leqslant G<\infty$ for any $\theta\in\Theta$ almost-surely. (iii) For any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|\leqslant D_{\mathcal{Z}}<\infty$ almost-surely.
It is common in online learning to assume boundedness of the parameter space, model gradient and prediction space (Boyd, 2004; Bishop and Nasrabadi, 2006). Note that in the well-specified setting, i.e., $\bar{g}_{t}=g(\bar{\theta}_{t},\cdot)$ for some $\bar{\theta}_{t}$, these two former conditions automatically imply the latter. We emphasize however that we do not assume this realizability condition, in contrast to most of the literature on decision-focused learning (Bennouna et al., 2024).
Without further assumptions on the sequence of costs $\bar{g}_{t}$ and models $g$, the problem can still be made arbitrarily hard. Therefore, we make an assumption coming from the classification (Mammen and Tsybakov, 1999; Tsybakov, 2004) and bandit literature (Zeevi and Goldenshluger, 2009; Perchet and Rigollet, 2013) about the margins of the cost function. Recall that we denote by $\{v_{i}\}_{i=1}^{K}$ the vertices of $\mathcal{W}$. We define for any $x\in\mathcal{X}$ and $i\in[K]$, $u_{i}(\theta,x)=\mathopen{}\mathclose{{\left\langle g(\theta,x),v_{i}}}\right\rangle$.
@@
There exist $C_{0}>0$ and $\beta\in(0,1]$, such that almost surely for any $t\in[T]$, $\theta\in\Theta$ and $\varepsilon\in\mathopen{}\mathclose{{\left[0,1}}\right]$,
@@
3 Online algorithms for decision-focused learning